OpenStudy (anonymous):

Hey, um i can't solve this math question can someone help me? If 8-x=4, and 10+y=12, then x-y= (A) 2 (B) 3 (C) 4 (D) 5 Don't give me the answer but help me figure out how to solve it. Thanks!

OpenStudy (hoblos):

8-x=4 then x=4 10+y=12 then y=2 so x-y= 4-2=2

OpenStudy (anonymous):

answer is A because 8-x=4 so u subtract the 8 and get -x=-4 and x=4. then for 10+x=12 subtract 10 and you get y=2..then subtract 4-2 and you get 2. so answer is A

OpenStudy (anonymous):

Thank you both!

OpenStudy (anonymous):

yea

OpenStudy (hoblos):

pleasured to help :D:D

OpenStudy (anonymous):

Hoblos can you look at my trig question please :(

OpenStudy (hoblos):

where is it ?

OpenStudy (anonymous):

Check my wall